Divide Rs. 140 between A and B so that A gets twice the money that B gets and Rs.20 more
Answer by macston(5194)

A=A's amount, B=B's amount
A+B= Rs. 140
A=2B+ Rs. 20 Substitute for A in first equation
2B + Rs. 20 +B=Rs. 140 Subtract Rs 20 from each side
3B=Rs. 120 Divide each side by 3
B=Rs. 40 ANSWER 1: B gets Rs. 40.
A=2(B)+ Rs. 20=2(Rs. 40)+Rs. 20=Rs. 80 + Rs. 20= Rs. 100
ANSWER 2: A gets Rs. 100.